OpenStudy (anonymous):

Which of the following is the vertex of the quadratic equation y = 6x^2 – 24x + 32?

OpenStudy (anonymous):

\[-\frac{b}{2a}\] which \[a=6,b=-24\] you get \[-\frac{-24}{2\times 6}=2\]

OpenStudy (anonymous):

(–2, –8) (–2, 8) (2, –8) (2, 8) those are the options..